I don't think there's any evidence that the noise eater is bad. That would change the behavior of the relaxation oscillation which is at 1 MHz ?
After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I
Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010
It's diodes should be replaced, based on it's age and performance.
RIN and noise eater bad. I will get a quote on this job.
The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956
Rana stated yesterday that there will be a vacuum control update in the close future. Witnesses : Rich, Chris and Dave
Can you give me this in writing?
Air condition maintenance is happening. It should be done by 10am
We will update the X circuit DCC page with an accurate schematic and photo.
I've uploaded reasonably high-resolution photographs of the uPDH box for the X-end and Y-end on their respective wiki pages. I've uploaded two photos for each box, one of the circuit board (I checked that these photos are clear enough that we can zoom in and read off component values if necessary), and one of the box with the peripherals not integrated into the circuit board (i.e. the minicircuits mixer ZAD-8+ and the little Pomona box that is an LP filter for the output from the mixer). Since I pulled the boxes out, I thought it might not be a bad idea to measure the TFs of these Pomona boxes and make sure nothing weird is going on, I'll put up some plots later.
Rana and I discussed some things to look at earlier today:
I also did a quick check of the behaviour of the Servo Gain potentiometer by checking the resistance at various positions of the knob - we had suspected that the potentiometer may be logarithmic, but I found that it was in fact linear. I'll put up a plot of the gain as a function of the Servo Gain knob position soon,(plot added) along with results from the other checks.
While disassembling the setup at the X-end to get the PDH box out, I noticed that the signal from the LO is going to the mixer through a Pomona box (no such Pomona box is used at the Y-end). I opened it up and found that it contains just a pair of capacitors in parallel, so it's a phase shifter?. The LO signal also goes through an attenuator. The mixer in both boxes is a ZAD-8+, so why is this part of the setup different?
Both PDH boxes are not hooked up at the moment, I will restore the setups at both ends after running a few more checks on the boxes...
Yesterday, I uploaded some EAGLE schematic files and a LISO source file for the green PDH servo electronics to the 40m LISO git repository. In doing so, I realized that the DCC document for the X box (D1400293) was not updated at the end of the electronics work we did in Aug/Sep 2014. This is entirely my fault.
The Y box document (D1400294) is currently accurate.
The missing information is that, as I posted In ELOG 10457, I ended up destroying our original X box, and replaced it with a spare from the ATF. It was restuffed to match the Y end box pretty much exactly. We will update the X circuit DCC page with an accurate schematic and photo.
Gautam tells me that he and Rana were looking at the outdated schematic and thinking about improvements, but at least some of this was already done back in 2014 (specifically, the resistors used to specify the AD8336 preamp gain were changed).
Summary of the work done today:
Alignment and other work on PSL table
As mentioned in a previous elog, the beatnote amplitude I obtained was tiny - so I checked the alignment of the two beams onto the PD. I did this as follows:
After doing all of this, I found a beatnote at ~-10dBm at a temperature of 45.3002 degrees on the Lightwave. The DC level was ~8V (~4V contribution from each beam).
PLL and frequency nosie measurements:
Pretty much the same procedure as that described in this elog was followed for setting up the PLL and taking the measurements, except that this time, I used the two SR560s in a better way to measure the open loop TF of the PLL. This measurement suggested a UGF of ~ 10kHz, which seems reasonable to me. I turned the 11MHz marconi off because some extra peaks were showing up in the beat signal spectrum. I judged that the beatnote was not large enough to require the use of an attenuator between the PD and the mixer. I was able to lock the PLL easily enough, and I've attached spectra of the control signal (both uncalibrated and calibrated). To calibrate the spectrum, I did a quick check to determine the actuator gain of the spare Lightwave laser, by sweeping the fast PZT with a low frequency (0.5Hz) 1Vpp sine wave, and looking at the peak in the beat signal spectrum move on the network analyzer. This admittedly rough calibration suggests that the coefficient is ~5MHz/V, consistent with the other Lightwave. Eric suggested a more accurate way to do this would be to match up spectra taken using this method and by locking the PLL by actuating on the FM input of the Marconi - I didn't try this, but given the relatively large low-frequency drifts of the beatnote that I was seeing, and that the control signal was regularly hitting ~2V (i.e shifting the frequency by ~10MHz), I don't think this is viable with a low MHz/V coefficient on the Marconi, which we found is desirable as described here.
The spare Lightwave frequency noise seems comparable to the other two measurements (see attachment #2). If anything, it is a factor of a few worse, though this could be due to an error in the calibration? I'm also not sure why the shapes of the spectra from today's measurement differ qualitatively from those in elog 11929 above ~7kHz.
Some random notes:
I didn't really appreciate this measurement until just now. IF you can save the DTT .xml file with all the traces in it (i.e. NOT just the plots), we should save this data for comparison plotting later. Perhaps Gautam can post the gzipped xml file for you into the log.
The accelerometers don't read any real noise below ~3 Hz, so we can't judge the difference down low, but this seems like a good measurement in the 5 - 100 Hz band.
Unfortunately I had closed all the DTT windows that Steve had used for the earlier plots. So I took the spectra again - there may be minor differences given that this measurement was taken at ~11pm at night. Anyways, plots and the xml data file are attached.
After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...
Objective: measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in
the larger optical table
The accelerometrs labeled as MC1 ( just north east of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:
MC1 to the floor at the north west leg of optical table.
MC2 is in the north east corner of the optical table
Atm2 was taken after table leg bolts were tighed at 40 ft/lb
The spectrum looks similar to ETMY except the Z direction
Conclusion: up to 20 Hz this set up is good.
I attempted to measure the frequency noise of the extra Lightwave NPRO we have that is currently sitting on the PSL table. I did the following:
I've turned the Lightwave NPRO back to standby for now, in anticipation of further trials later today. I've also restored the IMC.
The spectrum looks similar to ETMY (Guralp for below 3 Hz ) except the Z direction
The noisy fan belt on the roof of the control room was finally replaced on Friday, Jan 22 2016
PRM suspension damping restored after 4.1 Mag Ludlow earthquake.
The PDA photodetectors are DC coupled, so you cannot use them to go directly into the analyzer. Must use the DC block so that you can reduce the input attenuation on the B channel and then lower the drive amplitude.
Good policy for TF measurements: drive as softly as you can and still measure in a reasonable amount of time, but no softer than that.
There were a number of directories in /users/OLD/mott/PZT/2NPRO, I've used the data in Innolight_AM_New. Also, I am unsure as to what their "calibration" factor is to convert the measured data into RIN, so I've just used a value of 0.8, with which I got the plot to match up as close as possible to the plot in this elog. I also redid the measurement today, given that the laser parameters have changed. The main difference was that I used an excitation amplitude of +15dBm, and an "IF Bandwidth" of 30Hz in the parameter files for making these measurements, which I chose to match the parameters Mott used. There does seem to be a shift in some of the features, but the <100kHz area seems similar to the old measurement now.
Having put the PD back in, I also took measurements of the RIN with the input to the laser PZT terminated. There is no difference with the Noise Eater On or OFF!
Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region.
It looks like some of the features may have shifted in frequency. The previous measurement results can be found in /users/OLD/mott/PZT/2NPRO, can you plot the two AM measurements together?
I was trying to characterize the AM/PM response of the X end laser. I tried to measure the AM response first, as follows:
I hooked up the ALSX DFD output to the fibox, and used the adjustable delay line to set the phase properly. I recorded the noise on pianosa, and have attached it. Of course, this doesn't really capture the low frequency behavior.
Unrelated to this: I found the MC WFS turned off, and the loops ran away when turning them on. I tweaked the alignment, and reset the WFS offsets. Seems stable for now.
Is the black ref spectrum from this year or from May of 2015 or ?
I wonder if the noise is a bunch of fast spikes or if its a true broadband rumble. Maybe we can tell by looking at the analog DFD or PLL outputs?
That's a good news. Only quantitative analysis will tell us if it is true or not.
Also we still want to analyze the traffic with the new switch.
On a brighter note, I've only noticed one brief EPICS freeze all night. In addition, the wall StripTools seem totally contiuous since ~4pm, whereas I'm used to seeing some blocky shapes particularly in the seismic rainbow. Could this possibly mean that the old WiFi router was somehow involved in all this?
Earlier today, we did a bunch of stuff to see if we could improve the situation with the excess ALS-X noise. Long story short, here are the parameters that were changed, and their initial and final values:
X-end laser diode temperature: 28.5 degrees ---> 31.3 degrees
X-end laser diode current: 1.900 A ---> 1.942 A
X-end laser crystal temperature: 47.43 degrees ---> 42.6 degrees
PSL crystal temperature: 33.43 degrees ---> 29.41 degrees
PSL Diode A temperature: 21.52 degrees ---> 20.75 degrees
PSL Diode B temperature: 22.04 degrees ---> 21.3 degrees
The Y-end laser temperature has not yet been adjusted - this will have to be done to find the Y-beatnote.
Unfortunately, this does not seem to have fixed the problem - I was able to find the beatnote, with amplitude on the network analyzer in the control room consistent with what we've been seeing over the last few days, but as is clear from Attachment 1, the problem persists...
Some details not directly related to this work:
ETMX suspension damping restored.
We gave DRFPMI locking a shot, with the ALS out-of-loop noises as attached. I figured the ALSX noise might be tolerable.
After the usual alignment pains, we got to DRMI holding while buzzing around resonance. Recall that we have not locked since Koji's repair of the LO levels in the IMC loop, so the proper AO gains are a little up in the air right now. There were hopeful indications of arm powers stabilizing, but we were not able to make it stick yet. This is perhaps consistent with the ALSX noise making things harder, but not neccesarily impossible; we assuredly still want to fix the current situation but perhaps we can still lock.
While carrying out my end-table power investigations, I decided to take a quick look at the out-of-loop ALSX noise - see the attached plot. The feature at ~1kHz seems less prominent (factor of 2?) now, though its still present, and the overall noise above a few tens of Hz is still much higher than the reference. The green transmission was maximized to ~0.19 before this spectrum was taken.
We managed to access the trends for the green reflected and transmitted powers from a couple of months back when things were in their nominal state - see Attachment #2 for the situation then. For the X arm, the green reflected power has gone down from ~1300 counts (November 2015) to ~600 counts (january 2016) when locked to the arm and alignment is optimized. The corresponding numbers for the green transmitted powers (PSL + End Laser) are 0.47 (November 2015) and ~0.18 (January 2016). This seems to be a pretty dramatic change over just two months. For the Y-arm, the numbers are: ~3500 counts (Green REFL, Nov 2015), ~3500 counts (Green REFL, Jan 2016) ~1.3 (Green Trans, Nov 2015), ~1 (Green Trans, Jan 2016). So it definitely looks like something has changed dramatically with the X-end setup, while the Y-end seems consistent with what we had a couple of months ago...
Last week, Eric and I noticed that the green transmission levels at the PSL table seem much lower now than they did a month or two ago. To investigate this, I attempted to reproduce a power budget for the X endtable setup - see the attached figure (IR powers measured with calorimeter, green powers measured with Ophir power meter). A summary of my observations:
It is worth noting that two years ago, the IR power from the AUX-Y laser was ~280 mW, so we should still be getting "enough" green power for ALS?
Just an other local earthquake 3.6 Mag Ludlow, Ca
No obvious sign of damage
Just not just pedagogical ! Freq domain MISO coherence based subtraction estimation is much faster than calculating MISO WF. And since each bin is independent of each other, this gives us an estimate of how low the noise can go, whereas the Wiener filter is limited by Kramers-Kronig. We should be able to use this on the L1 DARM channel to do the noise hunting as well as estimating the subtraction efficacy of the pseudo channels that you and Rory come up with.
If you can code up a noise hunter example using DARM + a bunch of aux channels, we could implement it in the summary pages code.
The anticlimatic resolution to my subtraction confusion: Spectral leakage around 1Hz. Increasing the FFT length to 256 sec now shows that the FIR WF pretty much achieves the ideal subtraction.
If nothing else, it's good to have worked out how MISO coherence works.
No obvious sign of damage.
We checked the UGF of the AUX X PDH servo, found a ~6kHz UGF with ~45 degree phase margin, with the gain dial maxed out at 10.0. Laser current is at 1.90, direct IR output is ~300mW.
We recovered ALS readout of IR-locked arms. While the GTRX seemed low, after touching up the beam alignment, the DFD was reporting a healthy amount of signal. ALSY was perfectly nominal.
ALSX was a good deal higher than usual. Furthermore, there's a weird shape around ~1kHz that I can't explain at this point. It's present in both the IR and green beats. I don't suspect the DFD electronics, because the Y beat came through fine. The peak has moderate coherence with the AUX X PDH error signal (0.5 or so), but the shape of the PDH error signal is mostly smooth in the band in which the phase tracker output is wonky, but a hint of the bump is present.
Turning the PDH loop gain down increases the power spectrum of the error signal, obviously, but also smoothens out the phase tracker output. The PDH error signal spectrum in the G=10 case via DTT is drowning in ADC noise a bit, so we grabbed it's spectrum with the SR785 (attachment #2, ASD in V/rtHz), to show the smoothness thereof.
Finally, we took the X PDH box to the Y end to see how ALSY would perform, to see if the box was to blame. Right off the bat, when examining the spectrum of error signal with the X box, we see many large peaks in the tens of kHz, which are not present at the same gain with the Y PDH box. Some opamp oscillation shenanigans may be afoot... BUUUUUT: when swapping the Y PDH box into the X PDH setup, the ~1kHz bump is identical. ugh
In preparation for tonight's work, I did the following:
On the PSL table:
At the IOO Rack area:
At the X-end:
At the Y-end:
Having done all this, I checked the green transmission levels for both arms (PSL green shutter closed, after running ASS to maximize IR transmission). GTRY is close to what I remember (~0.40) while the best I could get GTRX to is ~0.12 (I seem to remember it being almost double this value - maybe the alignment onto the beat PD has to be improved?). Also, the amplitudes of the beatnotes on the network analyzer are ~-50dBm, and I seem to remember it being more like -25dBm, so maybe the alignment on the PD is the issue? I will investigate further in the evening. It remains to measure the OLTF of the X-end PDH as well.
I took several measurements today using the revised PLL scheme of using the Marconi just as an LO, and actuating on the Laser PZT to keep the PLL locked (I will put up a sketch soon). On the evidence of the attached plots (spectra of PLL control signal), I guess we can conclude the following:
Attachment #2: Measured OLG of PLL for the PSL+X and PSL+Y combinations. The UGF in both cases looks to be above 100 kHz, so I didn't do any calibration for the spectra attached. The gain on the SR560 was set to 200 for all measurements.
Attachment #3: Measured spectra of PLL control signal for various diode currents, with one reading from the PSL+Y combination plotted for comparison. When we took some data last night, Eric noted that there was a factor of ~6 increase in the overall frequency spectrum level at higher currents, I will update the plots with last night's data as well shortly. I found it hardest to keep the PLL locked at a diode current of 2.00 A across all measurements.
Attachment #4: Measured spectra of PLL control signal at two different crystal temperatures. There does not seem to be any significant dependance on temperature, although I did only do the measurement at two temperatures.
Attachment #4 Attachment #1: All the data used to make these plots (plus some that have yet to be added to the plots, I will update them).
Unrelated to this work:
When I came in this afternoon, I noticed that the PMC was unlocked. The usual procedure of turning the servo gain to -10dB and playing around with the DC output adjust slider on the MEDM screen did not work. Eric toggled a few buttons on the MEDM screen after which we were able to relock the PMC using the DC output adjust slider.
PMC locked and ETMX suspension damping restored.
Gautam will soon follow up with detailed analysis, but here is a brief summary of some of our activities and findings.
Please note that there is a long BNC cable still laid out from the IOO rack area to the X end table; watch your step!
EDIT 01/12/2016 6PM: I've updated the plots of the in-loop spectra such that they are calibrated throughout the entire domain now. I did so by inferring the closed-loop transfer function (G/(1-G)) from the measured open-loop transfer function (G), and then fitting the inferred TF using vectfit4 (2 poles). The spectra were calibrated by multiplying the measured spectra by the magnitude of the fitted analytic TF at the frequency of interest.
EricQ brought back one of the Marconis that was borrowed by the Cryo lab to the 40m today (it is a 2023B - the Marconi used for all previous measurements in this thread was 2023A). Koji had suggested investigating the frequency noise injected into the PLL by the Marconi, and I spent some time investigating this today. We tried to mimic the measurement setup used for the earlier measurements as closely as possible. One Marconi was used as a signal source, the other as the LO for the PLL loop. All measurements were done with the carrier on the signal Marconi set to 310MHz (since all our previous measurements were done around this value). We synced the two Marconis by means of the "Frequency Standard" BNC connector on the rear panel (having selected the appropriate In/Out configurations digitally first). Two combinations were investigated - with either Marconi as LO and signal source. For each combination, I adjusted the FM gain on the Marconi (D in the plot legends) and the overall control gain on the SR560 (G in the plot legends) such that their product remained approximately constant. I measured the PLL OLG at each pair to make sure the loop shape was the same throughout all trials. Here are the descriptions of the attached plots:
Attachment #1: 2023A as LO, 2023B as source, measured OLGs
Measured OLG for the various combinations of FM gain and SR560 gain tested. The UGF is approximately 30kHz for all combinations - the exceptions being D 1.6MHz, G=1e4 and D=3.2MHz, G=1e4. I took the latter two measurements just because these end up being the limiting values of D for different carrier frequencies on the Marconi.
Attachment #2: 2023A as LO, 2023B as source, measured spectra of control signal (uncalibrated above 30kHz)
I took the spectra down to 2Hz, in two ranges, and these are the stitched versions.
Attachment #3: 2023B as LO, 2023A as source, measured OLGs
Attachment #4: 2023B as LO, 2023A as source, measured spectra of control signal (uncalibrated above 30kHz)
So it appears that there is some difference between the two Marconis? Also, if the frequency noise ASD-frequency product is 10^4 for a healthy NPRO, these plots suggest that we should perhaps operate at a lower value of D than the 3.2MHz/V we have been using thus far?
As a quick trial, I also took quick spectra of the PLL control signals for the PSL+Aux X and PSL+Aux Y beat signals, with the 2023B as the LO (Attachment #5). The other difference is that I have plotted the spectrum down to 1 Hz (they are uncalibrated above 30Hz). The PSL+Y combination actually looks like what I would expect for an NPRO (for example, see page 2 of the datasheet of the Innolight Mephisto) particularly at lower frequencies - not sure what to make of the PSL+X combination. Also, I noticed that the amplitude of the PSL+Y beatnote was going through some large-amplitude (beat-note fluctuates between -8dBm and -20dBm) but low frequency (period ~10mins) oscillations. This has been observed before, not sure why its happening though.
More investigations to be done later tonight.
With the Y end laser, I was able to lock the PLL with a lower actuation range (1.6MHz/V), and with the PSL in both the free-running and MCL locked configurations.
I took spectra (attached) with the same actuation range (3.2 MHz/V) for the AUX X+PSL and AUX Y+PSL combinations (PSL shutter closed) just to keep things consistent. It looks like there is hardly any difference between the two combinations - could the apparent factor of 3 worse performance of the X end laser have been due to different actuation ranges on the Marconi?
I've not managed to take a spectrum for the proposed replacement Lightwave laser on the PSL table, though with Eric's help, I've managed to find the beatnote (at a temperature of 53.0195 degrees). I had to do some minor alignment tweaking for this purpose on the PSL table - the only optics I touched were the ones in the pink beam path in attachments 1 and 2 in this elog (the setup used to make the measurement is also qualitatively similar to attachment 3 in the same elog, except for the fact that we are feeding back to the Marconi and not the laser - a detailed sketch with specific components used will be put up later). I'll try and measure the frequency noise of this laser as well over the weekend and put up some spectra.
With regards to possibly switching out the Lightwave on the PSL table for the (faulty?) Innolight at the X end, I've verified the following:
It remains to characterize the beam coming out from the Lightwave laser and do a mode matching calculation to see if we can use the same optics currently in place (with slight rearrangement) to realize a satisfactory mode-matching solution - I've obtained a beam profiler to do this from Liyuan and have the software setup, but have yet to do the beam scan - the plan is to do this on the SP table, but we've put off moving the Lightwave laser off the PSL table until we (i) establish conclusively that the X end laser is malfunctioning and (ii) check the frequency nosie of the Lightwave relative to the Aux lasers currently at the ends.
The area around the Marconi is in a little disarray at the moment with a bunch of cables, SR560s, analyzers etc - I didn't want to disconnect the measurement setup till we're done with it. I have however turned both IR beat PDs on the PSL table off, and have reconnected the Marconi output to the Frequency Generation Unit and have set the carrier back to 11.066209MHz, +13dBm.
I configured a new wifi bridge for a GPIB Instruments.
The some facts are described on https://wiki-40m.ligo.caltech.edu/Network
The setting up wasn't so straight forward. I added more details there as a linked page.
One thing I had to do with the martian wifi router was that I had to separate the name of SSIDs for 2GHz and 5GHz networks.
Now the data download from Agilent is super fast!
The first establishing the connection takes the most of the time, and the data transfer takes literary nothing.
controls@pianosa|netgpibdata > time ./netgpibdata -i 192.168.113.167 -d AG4395A -a 10 -f meas01
Connecting to host 192.168.113.167, GPIB 10...
Data will be written into meas01.dat.
Parameters will be written into meas01.par.
Writing measurement data to file...
Writing to the parameter file.
Unless this is the limit from the way you guys set up the PLL, it seems like there's no difference between the two lasers that's of any import. So then the locking problem has been something else all along - perhaps its noise in the X-PDF lock somehow? PDH box oscillations?
Here are some results from measuring the PSL / AUX Y beat.
With the Y end laser, I was able to lock the PLL with a lower actuation range (1.6MHz/V), and with the PSL in both the free-running and MCL locked configurations. (In the latter, I had to do a bit of human-turning-knob servo to keep the control signal from running away). I also took a spectrum with the marconi detuned from the beat frequency, to estimate the noise from the PD+mixer+SR560.
It looks like the AUX X laser is about 3 times noisier than the Y, though the Y laser looks more like a 10^5 noise-frequency product, whereas I thought we needed 10^4.
Gautam is investigating the PSL / AUX PSL beat with Koji's setup now.
The next step is to compare this data with the same measurement with the PSL and the AUX laser on the PSL table (or the end Y laser). If these show a lot lower noise level, we can say 1) the x-end laser is malfunctioning and 2) the y-end and AUX laser on the PSL are well low noise.
Here is some of the promised data. As mentioned, changing diode current and crystal temperature didn't have much effect on the frequency noise spectrum; but the spectrum itself does seem too high for our needs.
At each temperature, we started measuring the spectrum at 1.8A, and stepped the current up, hoping to reach 2.0 A.
At 47.5 C, we were able to scan the current from 1.8 to 2.0 A without much problem. At 49.0C, the laser mode would hop away above 1.95A. At 50.4C it would hop away above 1.85A. The spectra were not seen to change when physically disconnecting the PZT actuation BNC from the rear of the laser.
The flattening out at the upper end is likely due to the SR560 output noise. I foolishly neglected to record the output spectrum of it, but with the marconi external modulation set to 3.2MHz/V, the few Hz/rtHz above 20k translates to a signal on the order of uV/rtHz, which seems reasonable.
Data and code attached.
The puzzle continues...
I found some reference for computing "multicoherence," which should properly estimate the potential MISO subtraction potential in situations where the witness channels themselves have nontrivial coherence. Specifically, I followed the derivations in LIGO-P990002. The underlying math is related to principal component analysis (PCA) or gram-schmidt orthogonalization.
This produced the following results, wherein the Wiener subtraction is still below what the coherences predict.
I've attached the data and code that produced this plot.
Brief summary of tonights work:
Our "requirement" for the end laser is as follows: We expect to (and have in the past) achieved ALS sensitivity of 1Hz/rtHz at 100 Hz. If the end PDH loop is 1/f from 100Hz-10kHz, then we have 40dB of supression at 100Hz, meaning the free running AUX laser noise should be no more than 100Hz/rtHz at 100Hz.
So, if we expect both the PSL and AUX lasers to have this performance when free running, we would get the green curve below. We do not.
I'll post more details about the exact currents, temperatures and include calibrated plots for the >30kHz range later. Here's the OLG for kicks.
The new wifi router, a Netgear R6400, has been installed, next to the old one which is disconnected (but not yet removed).
Same SSID, and I've added only the wireless MAC addresses of viviana, paoloa and asia, the three thinkpads inside.
Qualitatively, dataviewer at the X end seems pretty snappy. I'll do some more quantitative comparison of the two routers at some point soon. I will update the wiki, too.
Projector light bulb ordered
[ Manasa, Ericq and Steve ]
Vivitek D952HD with 186 hours installed.
IFO restored after 4.5 Mag Banning, Ca earthquake.
In the modern times, people use glue traps to catch rats instead of springs. They are less hazardous to people and don't spread rat fluid on the floor.